find the equation (in any correct form) of a straight line through (-5,2) which is perpendicular to the line whose equation is 4x+7y=9
Answer by scott8148(6628)

to find a perpendicular line; swap the x and y coefficients and change one of the signs
___ substitute the given point to find the new constant term

7x - 4y = ? ___ 7(-5) - 4(2) = -43

7x - 4y = -43